A mixture of 6% disinfectant solution is to be made from 7% and 3% disinfectant solutions. How much of each solution should be used if 36 gallons of 6% solution are needed?
Answer by Alan3354(69443)

A mixture of 6% disinfectant solution is to be made from 7% and 3% disinfectant solutions. How much of each solution should be used if 36 gallons of 6% solution are needed?
----------
t = amount of 3%
s = amount of 7%
----
t + s = 36 (total solution)
3t + 7s = 6*36 (total amount of disinfectant)